# Can repeating and non-repeating FRBs be drawn from the same population?

### v1

@@ Appendix B Correlation between rate and maximal energy
$\frac{N_{\geq 1}}{N_{\geq 0}}\!\approx\!\left(\frac{E_{\rm lim}}{E_{*}}\right)^{-\frac{a}{b}}\left\{\begin{array}[]{ll}\!\lambda_{\rm lim}^{a}&0\!<\!a\!<\!1\\ \!\lambda_{\rm lim}&a>1\ .\end{array}\right.\approx\lambda_{\rm 0,*}\left\{\begin{array}[]{ll}\!\left(\frac{E_{\rm lim}}{E_{*}}\right)^{1-\gamma}&0\!<\!a\!<\!1\\ \!\left(\frac{E_{\rm lim}}{E_{*}}\right)^{\frac{1-a}{b}+1-\gamma}&a>1\ .\end{array}\right.$ (B5)
@@
For $b>0$, sources dominating the CHIME sample, satisfy $\lambda_{\rm CH}(E_{\rm c})=\lambda_{0,*}(E_{*}/10^{26}\mbox{erg Hz}^{-1})^{1-\gamma}(E_{\rm c}/E_{\rm*})^{1/b+1-\gamma}\approx 1.3\times 10^{-5}(E_{\rm c}/E_{\rm lim})^{1/b+1-\gamma}$ with $E_{\rm c}\geq E_{\rm lim}$, such that the results depend on $b$ for $E_{\rm c}>E_{\rm lim}$. This is more than three orders of magnitude greater than the rate of bursts per SGR as extrapolated to $E_{\rm lim}$. We see also that the (isotropic equivalent) energy needed to power this activity is $E_{\rm rel}\approx t\mathcal{R}_{\rm CH}(E_{\rm c})E_{\rm c}\approx 1.6\times 10^{40}(t/30\mbox{ yr})(E_{\rm c}/E_{\rm lim})^{1/b+2-\gamma}\mbox{ erg}$, where $t$ is the time over which the source has been active at this level. This energy is many orders of magnitude lower than the magnetic energy reservoir of a magnetar ($E_{\rm mag}=1.6\times 10^{49}(B/10^{16})^{2}\mbox{ erg}$), even after accounting for the fact that typically only a small fraction of the dissipated energy may be channeled towards the radio band. Since we infer that $a-1\ll 1$, a wide range of intrinsic source repetition rates should be probed by CHIME sources. Indeed Beniamini & Kumar (2025) found that for R1, the energy requirements result in $E_{\rm rel}\approx 8.4\times 10^{44}\mbox{ erg}$ for the same beaming factor, radio efficiency and activity time, about $3\times 10^{5}$ times larger than for the typical CHIME source, as calculated above.

### v2

@@ Appendix C Correlation between rate and maximal energy
$\frac{N_{\geq 1}}{N_{\geq 0}}\!\approx\!\left(\frac{E_{\rm lim}}{E_{*}}\right)^{-\frac{a}{b}}\left\{\begin{array}[]{ll}\!\lambda_{\rm lim}^{a}&0\!<\!a\!<\!1\\ \!\lambda_{\rm lim}&a>1\ .\end{array}\right.\approx\lambda_{\rm 0,*}\left\{\begin{array}[]{ll}\!\left(\frac{E_{\rm lim}}{E_{*}}\right)^{1-\gamma}&0\!<\!a\!<\!1\\ \!\left(\frac{E_{\rm lim}}{E_{*}}\right)^{\frac{1-a}{b}+1-\gamma}&a>1\ .\end{array}\right$. (C5)
@@
For $b>0$, sources dominating the CHIME sample, satisfy $\lambda_{\rm CH}(E_{\rm c})=\lambda_{0,*}(E_{*}/10^{26}\mbox{erg Hz}^{-1})^{1-\gamma}(E_{\rm c}/E_{\rm*})^{1/b+1-\gamma}\approx 1.3\times 10^{-5}(E_{\rm c}/E_{\rm lim})^{1/b+1-\gamma}$ with $E_{\rm c}\geq E_{\rm lim}$, such that the results depend on $b$ for $E_{\rm c}>E_{\rm lim}$. This is more than three orders of magnitude greater than the rate of bursts per SGR as extrapolated to $E_{\rm lim}$. We see also that the (isotropic equivalent) energy needed to power this activity is $E_{\rm rel}\approx t\mathcal{R}_{\rm CH}(E_{\rm c})E_{\rm c}\approx 1.6\times 10^{40}(t/30\mbox{ yr})(E_{\rm c}/E_{\rm lim})^{1/b+2-\gamma}\mbox{ erg}$, where $t$ is the time over which the source has been active at this level. This energy is many orders of magnitude lower than the magnetic energy reservoir of a magnetar ($E_{\rm mag}=1.6\times 10^{49}(B/10^{16})^{2}\mbox{ erg}$), even after accounting for the fact that typically only a small fraction of the dissipated energy may be channeled towards the radio band. Since we infer that $a-1\ll 1$, a wide range of intrinsic source repetition rates should be probed by CHIME sources. Indeed Beniamini & Kumar (2025) found that for FRB 20121102A, the energy requirements result in $E_{\rm rel}\approx 8.4\times 10^{44}\mbox{ erg}$ for the same beaming factor, radio efficiency and activity time, about $3\times 10^{5}$ times larger than for the typical CHIME source, as calculated above.